### v1

# Cavity Control of Topological Qubits: Fusion Rule, Anyon Braiding and Majorana-Schrödinger Cat States

Abstract
We investigate the impact of introducing a local cavity within the center of a topological chain, revealing profound effects on the system’s quantum states. Notably, the cavity induces a scissor-like effect that bisects the chain, liberating Majorana zero modes (MZMs) within the bulk. Our results demonstrate that this setup enables the observation of non-trivial fusion rules and braiding –key signatures of non-Abelian anyons– facilitated by the spatially selective ultra-strong coupling of the cavity photon field. These MZM characteristics can be directly probed through fermionic parity readouts and photon Berry phases, respectively. Furthermore, by leveraging the symmetry properties of fermion modes within a two-site cavity, we propose a novel method for generating MZM-polariton Schrödinger cat states. Our findings present a significant advancement in the control of topological quantum systems, offering new avenues for both fundamental research and potential quantum computing applications.
Introduction.— Topological quantum computing is a holy grail of current research on quantum technologies [1]. Its key building blocks are non-abelian anyon quasi-particles in topological matter [2, 3, 4]. Clear experimental evidence on such anyons is still missing but recent advances in precision control of hybrid semiconductor-superconductor nanowires and quantum dot arrays have brought researchers closer to their realization [5, 6]. This opens new possibilities in designing novel experimental frameworks while making conceptual questions of great relevance.
The essential physics of 1D Ising anyons is captured by the Kitaev chain (KC) model [7]. Each local Dirac fermion of the chain can be expressed in terms of a pair of Majorana fermions (MF). A key consequence is the existence of a topological phase in which a zero energy mode exists. It consists of two MF each one localized at one edge of the chain. Occupied and unoccupied options of such state have zero energy, forming a non-local qubit. Due to the splitting of the Majorana zero mode (MZM), quantum information is stored in both halves (data redundancy). If one half of the mode suffers interference, there is still enough information stored in the other half to proceed with the computation. Moreover, MZM become decoupled from the rest of the KC being an excellent candidate to present non-abelian statistics. Topological protection is produced through braiding and measurement of the non-abelian anyon character of MZM.
Recently, braiding and quantum teleportation of delocalized Majorana qubits have been established in KC manipulated by superconducting resonators [8, 9]. This remarkable experimental achievement prompts us to elaborate on extended scenarios to theoretically investigate for further non-local features in condensed matter systems aiming to put on firm grounds topological quantum computation routines. The most popular scheme proposed to achieve anyon braiding in topological 1D wire networks relies on a T-junction geometry [10, 11, 12]. The goal of most studies has been to explore how the pairing between different MF featuring an MZM evolves under a cyclic tuning of different coupling schemes. These simulations can be challenging both experimentally and theoretically. On the other side, the question of what might happen with an anyon fusion or braiding protocol relying on a purely geometric quantum state property has received far less attention.
The present work aims to pave the way to an alternative to multi-wire T-junction schemes: the analysis of the potential capabilities of a wire topological network controlled by the coupling to a bosonic environment in the regime of cavity quantum electrodynamics. This would open the use of photons and topological matter-connected states for new strategies to harness non-classical aspects of both matter and radiation ranging from fundamental physics aspects to potential applications in quantum information/computation issues. This becomes even more relevant in powerful platforms with novel quantum materials strongly coupled to optical and microwave cavities [13, 14, 15, 16, 17, 18] where photon quantum states, specifically coherent and squeezed states, show great promise in this regard due to the ease of availability and well-established control techniques.
In stark contrast to previous proposals that aimed the demonstration of Majorana fusion rule using diverse connecting elements (quantum dots or electrostatic gates), to mediate the exchange between different KC edge MZMs [11, 12, 19, 20], our proposed KC-cavity set up is based on a single KC in the deep topological phase. In this way, the cavity acts as a scissor, splitting the original KC into two halves and consequently generating two new MZMs (for a fully electrostatic cutting way see [21]). The basic idea of previous works with cavity-controlled MZM systems [22, 23, 24] was primarily focused on the lowest-order photon-MZM coupling effects, i.e. where pre-defined MZM already existed before embedding the topological matter within the cavity, instead of a genuine full polariton image. Here, we propose a significantly distinct physics framework: the local cavity acts as a topological scissor [25], giving rise to a set of MZMs at the ends of the two parts of the original KC outside the cavity. Furthermore, it enables the simulation of their exchange by gathering up the precise Berry phase of the cavity field under an adiabatic cyclic variation of the matter-light coupling strength. As we will demonstrate below, for low and intermediate coupling strengths the polariton state is an entangled matter-light state. Nevertheless, in the ultra-strong coupling (USC) regime, the asymptotic decoupling of light and matter degrees of freedom [26, 27] enables both anyon fusing and braiding processes, mediated by photons, to occur in a KC-cavity system. Moreover, highly non-classical photon states in the cavity, Schrödinger cat states assisted by Majorana nonlocal qubits, become accessible. Such features turn out to be within reach in the USC regime due to the burgeoning interest on new cavity quantum electrodynamics (cQED) architectures and material supports [14, 15, 16]. The USC is particularly interesting for going beyond such first-order effects. The transition between different photon coherent states goes with corresponding modifications of fermion correlations. Moreover, of particular interest is the fact that in the asymptotic decoupling regime, cavity decoherence effects turn out to be innocuous for the MZM dynamics.
Model system.— We consider a single KC with vanishing chemical potential $\mu=0$, and identical hopping ($t$) and pairing ($\Delta$) terms, i.e. $t=\Delta$, the so-called sweet-spot for a quantum dot realization of the minimal KC [21]. For vanishing chemical potential the Kitaev Hamiltonian takes a diagonal form in terms of new non-local fermion operators $\hat{d}_{j}$ and ${\hat{d}_{j}^{\dagger}}$, i.e. $\hat{H}_{\rm KC}=2\Delta\sum_{j=1}^{N-1}\left[\hat{d}_{j}^{\dagger}\hat{d}_{j}-\frac{1}{2}\right]=i\Delta\sum_{j=1}^{N-1}\hat{\gamma}_{j,2}\hat{\gamma}_{j+1,1}$ (see Supplemental Material (SM) [*[SeetheSupplementalMaterialat][fordetailsofthecalculationsandderivations.]SM_Quiroga] and Ref. [29] for their definition in terms of both original Kitaev local operators ($\hat{c}_{j}^{\dagger},\hat{c}_{j}$) or Majorana fermions ($\hat{\gamma}_{j,1},\hat{\gamma}_{j,2}$)). The single mode microcavity Hamiltonian takes the usual form $\hat{H}_{\rm C}=\omega\hat{a}^{\dagger}\hat{a}$, where $\omega$ is the photon quantum energy ($\hbar=1$). Finally, the KC-cavity coupling term is given by $\hat{H}_{X}=\frac{1}{\sqrt{n_{\rm Cav}}}\sum_{j\in{\rm Cav}}\lambda_{j}\hat{c}_{j}^{\dagger}\hat{c}_{j}\left(e^{-i\phi_{j}}\hat{a}^{\dagger}+\hat{a}e^{i\phi_{j}}\right)$ with $\lambda_{j}$ the coupling strength between the KC $j$-site and the cavity field, ${\rm Cav}$ is the set with the site indices interacting with the cavity and $\phi_{j}$ controllable phases [30].
[…]
$\hat{H}_{1}=-2\Delta\hat{\sigma}_{z}+\omega\hat{a}^{\dagger}\hat{a}+\frac{\lambda}{2}\left(e^{-i\phi}\hat{a}^{\dagger}+e^{i\phi}\hat{a}\right)\left(1-\hat{\sigma}_{x}\right).$ (1)
Figure 1: Schematic representation of Kitaev chain-cavity systems. In the topological phase, a Kitaev chain is schematically represented by colored balls and links. The balls represent Majorana fermions $\gamma_{j+1,1}$ and $\gamma_{j,2}$ with different colors, indicating the localization of a free MZM as a torus. The cavity is depicted by two blue mirrors, operating as natural topological scissors. The KC-cavity interaction in the USC regime allows the formation of local Dirac fermions located inside the cavity. (a) corresponds to the $1CK$ case while (b) depicts the $2CK$ case. The radiation-matter interaction is represented by $\lambda_{r}$, with $r=1,2$.
Similarly, for a cavity enclosing two contiguous KC sites (from now on $2CK$ case, Fig. 1(b), sites $j=s$ and $j=s+1$, or equivalently only three $d$-fermion operators ($\hat{d}_{s-1},\hat{d}_{s}$ and $\hat{d}_{s+1}$) are brought into play so just $4$ matter states are involved in the full system quantum evolution, a two-qubit effective Rabi-like Hamiltonian arises:
$\begin{split}\hat{H}_{2}&=-\Delta\left(\hat{\sigma}_{z,1}+\hat{\sigma}_{z,2}+\hat{\sigma}_{z,1}\hat{\sigma}_{z,2}\right)+\omega\hat{a}^{{\dagger}}\hat{a}\\ &+\frac{1}{2\sqrt{2}}\sum_{r=1}^{2}\lambda_{r}\left(e^{-i\phi_{r}}\hat{a}^{\dagger}+e^{i\phi_{r}}\hat{a}\right)\left(1-\hat{\sigma}_{x,r}\right),\end{split}$ (2)
where $\hat{\sigma}_{\nu,r}$, with $\nu=x,y,z$, denote Pauli matrices for spin $r=1,2$, complex valued $\lambda_{1}e^{-i\phi_{1}}$ and $\lambda_{2}e^{-i\phi_{2}}$ represent the coupling strengths at sites $s$ and $s+1$, respectively (see SM [28]). Restriction to the even fermion parity sector has been kept. Notice that for both $1CK$ and $2CK$ cases, the resonance condition between matter and photonic sub-systems is $\omega=4\Delta$ and the $d$-fermion parity is a constant of motion, i.e. a single photon can only produce changes on the $d$-fermion occupation number by $0,\pm 2$.
In the USC regime, the $2CK$ ground state displays a clear decoupling between matter and photon states for any pair ($\lambda_{1},\lambda_{2}$), as shown in Fig. 2, except for the special point $\lambda_{2}=-\lambda_{1}$ where a superposition of even and odd Schrödinger cat states occurs. Simultaneously, a sudden jump of the coherent state displacement $\alpha$ takes place when crossing this special point. Signatures of those transitions are also imprinted in the photon distribution (see Fig. 2), allowing for optical readout schemes. These special features will be exploited below to set up a braiding protocol.
Figure 2: $2CK$ ground state phase diagram in the USC regime. Inhomogeneous system, i.e. real valued $\lambda_{2}\neq\lambda_{1}$. Full numerical (open circles) and USC theoretical results (broken line) for the photonic coherent state displacement $\langle\hat{a}\rangle$ are put on top: $\lambda_{1}/2\Delta=4$ and resonance condition $\omega/2\Delta=2$. See main text for details.
Fusion rule.— MZMs are Ising anyons obeying the fusion rule $\sigma\times\sigma=1+\Psi$, where two Ising anyons ($\sigma$) fuse into either a vacuum ($1$) or a standard Dirac fermion ($\Psi$[21, 31]. Fusion rule experiments are in general simpler than braiding experiments, thus they are first-ever candidates for testing the non-Abelian features of MZMs.
In the present cavity setup, for testing the non-trivial fusion rule only a pair of MZMs is required. We consider the fusion rule in the $1CK$ case by adiabatically increasing the KC-cavity coupling strength. Thus, we start with a single uncoupled KC in its ground state featuring a pair of MZMs located at the chain edges. During the whole process, the total fermionic parity of the KC states is fixed to be even ($+1$) as we start with the ground state of a non-interacting KC-cavity, i.e. $|\chi_{1}\rangle=|-\rangle_{z}\otimes|0_{\rm ph}\rangle$ where $|0_{\rm ph}\rangle$ denotes the photon vacuum state. After the cavity is strongly coupled to the KC two new MZMs emerge at the neighbor sites of the cavity as previously demonstrated in Ref. [25].
One immediate consequence of the $H_{1}$ structure in Eq. (1) is the expression for its USC limit eigenstates: if $\lambda\gtrsim\omega\gg\Delta$, the matter-light eigenstates become uncoupled, given by the separable states $|\Psi\rangle\approx|-\rangle_{x}\otimes|\alpha\rangle$ where $\hat{\sigma}_{x}|-\rangle_{x}=-|-\rangle_{x}$ and the photon state is nothing but the coherent state $|\alpha\rangle$ with $\alpha=-\lambda/\omega$. Thus, the final matter state, after adiabatically turning on the KC-cavity coupling until the USC regime is reached, turns out to be $|-\rangle_{x}=\left(\lvert\circ\circ\rangle-|\bullet\bullet\rangle\right)/\sqrt{2}$ indicating once more that the cavity coupling doesn’t alter the total fermionic parity. Therefore, the final matter state is an equal probability superposition of states with no fermions on each side of the cavity and states with a single fermion on each side. The fusion outcome becomes probabilistic, not deterministic, as clearly indicated by the resulting KC-cavity state. This result corresponds to a positive test of the fusion protocol of two Ising anyons. In the case of quantum dot-based Kitaev systems the readout of the fusion outcome can be detected by standard capacitance measurements or tunneling signatures [32].
Figure 3: Fusion rule protocol in a $1CK$ system. In panel (a), we depict the ground state fermion correlation function $P(s-1,s+1)$, with the dashed blue line representing the resonance condition. In panel (b), for the resonant case, we consider a time-dependent linear ramping of the radiation-matter interaction coupling strength, $\lambda=vt=t/t_{a}$ with $v=t_{a}^{-1}$ the ramping rate, where the evolution time $t\in\left[0,4t_{a}\right]$.
We numerically test the validity of the fusion protocol in $1CK$ system by diabatically turning on the matter-light coupling strength. For that purpose we evaluate the fermion correlation function of the two nearest neighbor Majorana fermions to the cavity, $P(s-1,s+1)=i\langle\gamma_{s-1,2},\gamma_{s+1,1}\rangle=-\langle\hat{\sigma}_{x}\rangle$. Figure 3(a) shows a 3D plot of that correlation in the ground state as a function of photon energy and coupling strength ($\phi=0$). To test for diabatic errors, we show the same correlation function in Fig. 3(b) for time-dependent cases with different velocities of turning on of the matter-photon coupling strength (different ramping rates). Notably, we obtain an excellent agreement between analytical and numerical results, even well beyond the realm of applicability of the USC regime, for sufficiently slow ramping processes.
Braiding protocol.— As a further step, we propose a cavity-chain setup to perform an MZM braiding/exchange process. Most recipes for simulations of braiding protocols in 1D topological systems rely on either Floquet dynamics (Majorana time crystals [33]) or quantum dot mediated exchange [21, 34]. The former case demands extreme timing control while the latter scheme requires a vanishing dot energy level splitting which hampers an adiabatic evolution harming the goal of a geometric phase $\pi/4$, as required for a perfect exchange between two MZMs. Our goal is to reach the same $\pi/4$-phase by performing an adiabatic cyclic process on the cavity itself, i.e. on the coherent state $|\alpha\rangle$, by leveraging the separable property of the ground state (GS) in the USC regime, i.e. $\lvert{\rm GS}\rangle\simeq\lvert\chi\rangle\otimes|\alpha\rangle$ with $\lvert\chi\rangle$ a matter KC state. In our setup, the polariton KC-cavity state can be coupled to the Majorana state, and so is not involved in braiding, only initialization (and possibly readout).
Consider a $2CK$ acting on chain sites $j=s$ (coupling strength $\lambda_{1}e^{i\phi_{1}}$) and $j=s+1$ (coupling strength $\lambda_{2}e^{i\phi_{2}}$), governed by $H_{2}$ in Eq. (2). Let us start by fixing $\lambda_{1}>0$ in the USC regime and $\phi_{1}=0$, without any loss of generality, while $\lambda_{2}$ and $\phi_{2}$ can be tunable. The MF just to the left of the cavity, i.e. $\hat{\gamma}_{L}=\hat{\gamma}_{s-1,2}$, always behaves as a free MZM [25] while $\hat{\gamma}_{R}=\hat{\gamma}_{s+1,1}$ acquires the MZM condition only in the $\lambda_{2}\rightarrow 0$ regime. We will be focusing on the braiding process between these two MFs, $\hat{\gamma}_{L}$ and $\hat{\gamma}_{R}$.
For $\phi_{1}=0$, the composite ground state in the USC regime becomes
$\lvert{\rm GS}_{\pm}\rangle\simeq\begin{cases}-i\lvert\Psi_{+}\rangle\otimes|\alpha_{+}\rangle,&\text{for }\phi_{2}=0,\\ +i\lvert\Psi_{-}\rangle\otimes|\alpha_{-}\rangle,&\text{for }\phi_{2}=\pi.\end{cases}$ (3)
Where $\lvert\Psi_{\pm}\rangle=(\lvert\bullet_{L,R},\bullet_{c_{1}},\circ_{s+1}\rangle\pm\lvert\circ_{L,R},\bullet_{c_{1}},\bullet_{s+1}\rangle)/\sqrt{2}$, $\alpha_{+}=-(\lambda_{1}+\lambda_{2})/(\sqrt{2}\omega)$, and $\alpha_{-}=-\lambda_{1}/(\sqrt{2}\omega)$ regardless the value of $\lambda_{2}\leq\lambda_{1}$. In the 1CK limit, i.e. $\lambda_{2}\rightarrow 0$, the system’s ground state turns out to be
$\begin{split}\lvert{\rm GS}\rangle&\simeq\frac{1}{\sqrt{2}}\left(\lvert{\rm GS}_{+}\rangle-\lvert{\rm GS}_{-}\rangle\right),\\ &=\lvert\bullet_{L,R},\bullet_{c},\circ_{s+1}\rangle\otimes|\alpha\rangle.\end{split}$ (4)
With $\alpha=-\lambda_{1}/(\sqrt{2}\omega)$. The braiding or exchange process between $\hat{\gamma}_{L}$ and $\hat{\gamma}_{R}$ MFs, as described by the operator $\hat{U}_{B}^{(L,R)}=e^{\frac{\pi}{4}\hat{\gamma}_{L}\hat{\gamma}_{R}}=\frac{1}{\sqrt{2}}\left(1+\hat{\gamma}_{L}\hat{\gamma}_{R}\right)$ (performing the exchange transformations $\hat{\gamma}_{L}\rightarrow\hat{\gamma}_{R}$ and $\hat{\gamma}_{R}\rightarrow-\hat{\gamma}_{L}$), when acting on the ground state produces the result $\hat{U}_{B}^{(L,R)}\lvert{\rm GS}\rangle=e^{-i\frac{\pi}{4}}\lvert{\rm GS}\rangle$. By using a local phase shifter acting just on the $j=s+1$ chain site [35, 36, 37, 30], a controlled complex coupling strength $\lambda_{2}e^{i\phi_{2}}$ should be possible enabling a cyclic evolution of the cavity coherent state $|\alpha\rangle$ and consequently to access to a specific Berry phase of the whole matter-photon state (dynamic phases which depend on the system’s specific time evolution may be experimentally canceled out [38, 39]).
Figure 4: Signature of MZM Braiding in a 2KC Setup. In panel (a), we depict the numerical relationship between the couplings $\lambda_{2}$ as a function of $\lambda_{1}$, where the Berry phase is exactly $\pi/4$. In the inset of panel (a), we show the behavior of the Berry phase by numerical evaluation of Eq. (5) for several values of $\lambda_{1}$. In panel (b), we illustrate the landscape of parity between $\gamma_{L}$ and $\gamma_{R}$.
Now, we focus on recovering the same $e^{-i\frac{\pi}{4}}$ global phase factor on the GS by performing an adiabatic cyclic evolution of the coherent cavity field (Berry phase). The GS adiabatic evolution is fully contained in the two-dimensional sub-space spanned by the states $\{\lvert{\rm GS}_{+}\rangle,\lvert{\rm GS}_{-}\rangle\}$, as previously assured (for details see the SM). We perform a cyclic evolution of the GS by fixing a $\lambda_{2}>0$ value while varying the phase $\phi_{2}=\phi$ of the coupling strength, $0\leq\phi\leq 2\pi$. Thus, for the GS we have $\lvert{\rm GS}(\phi)\rangle=C_{+}(\phi)\lvert{\rm GS}_{+}(\phi)\rangle+C_{-}(\phi)\lvert{\rm GS}_{-}\rangle$. Notice that the state $\lvert{\rm GS}_{-}\rangle$ does not depend on the phase $\phi$. We proceed with an analytical calculation of the Berry phase using the sub-space-confined evolution of the GS. By direct calculation, we obtain:
$\varphi_{B}=\int_{0}^{2\pi}\frac{d\phi}{2\omega}\delta E_{\phi}\left[1-\frac{\delta E_{\phi}}{\sqrt{4\Delta^{2}\lvert\langle\alpha_{-}\lvert\alpha_{+}(\phi)\rangle\lvert^{2}-\delta E^{2}_{\phi}}}\right].$ (5)
Where, we are using $\delta E_{\phi}=\omega(|\alpha_{+}(\phi)|^{2}-|\alpha_{-}|^{2})$, $\alpha_{-}=-\lambda_{1}/(\sqrt{2}\omega)$, and $\alpha_{+}(\phi)=-(\lambda_{1}+\lambda_{2}e^{i\phi})/(\sqrt{2}\omega)$. If $\Delta\ll\lambda_{2}$, the Berry phase $\varphi_{B}\rightarrow\pi\lambda_{2}^{2}/\omega^{2}$, coinciding exactly with the Berry phase of a pure coherent state with $\alpha=-\lambda_{2}/(\sqrt{2}\omega)$ under an adiabatic cyclic evolution [40]. A numerical evaluation of Eq. (5) allows us to make a plot of $\lambda_{2}$ as a function of $\lambda_{1}$ to get a Berry phase of just $\pi/4$, as shown in Fig. 4(a). As it is evident from that figure to reach the appropriate Berry phase the coupling strength with the right KC site $\lambda_{2}\ll\lambda_{1}$, validating the ground state structure displayed by Eq. (4). Measurement of the photon Berry phase in superconducting cavity systems has been widely discussed previously [41].
The parity $P_{L,R}=i\langle\hat{\gamma}_{L}\hat{\gamma}_{R}\rangle$, between $\hat{\gamma}_{L}$ and $\hat{\gamma}_{R}$ MFs, is plotted in Fig. 4(b) as a function of $\lambda_{2}e^{i\phi_{2}}$, for $0\leq\lambda_{2}\leq\lambda_{1}$ and $0\leq\phi_{2}\leq 2\pi$. $P_{L,R}\rightarrow-1$ for $\lambda_{2}\approx 0$ in agreement with the GS given by Eq. (4). The solid red line shows the corresponding results for varying real $\lambda_{2}$ values from the positive to the negative region. At the same time, the closed curve depicts the $P_{L,R}$ evolution just at the $\lambda_{2}$ value producing the correct Berry phase of $\pi/4$, simulating the perfect braiding/exchange between $\hat{\gamma}_{L}$ and $\hat{\gamma}_{R}$ MFs.
Majorana-Schrödinger cat states.— Photon Schrödinger cat states have been intensively studied in circuit quantum electrodynamics scenarios since they offer the opportunity of investigating fundamental tests of quantum theory, for encoding logical qubits and for developing error correction codes [42]. A Schrödinger cat-like state can be written as $\lvert{\rm\bf Cat}_{\pm}\rangle=\left(\lvert\alpha\rangle\pm\lvert-\alpha\rangle\right)/N_{\pm}$ with normalization constant $N_{\pm}^{2}=2\left(1\pm\exp\left[-2|\alpha|^{2}\right]\right)$ and sign $+$ (--) for even (odd) cat state. Interestingly enough, in the USC regime, a $2CK$ system supports photon Schrödinger cat states coupled to Majorana fermion states at the special coupling values $\lambda_{1}=\lambda_{2}=\lambda$, $\phi_{1}=0$, and $\phi_{2}=\pi$. The $2CK$ ground state becomes:
$\begin{split}\lvert{\rm GS}\rangle&\simeq\lvert+,-\rangle_{x}\otimes|-\alpha\rangle+\lvert-,+\rangle_{x}\otimes|\alpha\rangle,\\ &\simeq\lvert\tilde{\Psi}_{+}\rangle\otimes\lvert{\rm\bf Cat}_{+}\rangle+\lvert\tilde{\Psi}_{-}\rangle\otimes\lvert{\rm\bf Cat}_{-}\rangle.\end{split}$ (6)
Where, we are using $\alpha=-\lambda/(\sqrt{2}\omega)$, $\lvert\tilde{\Psi}_{\pm}\rangle=(\lvert\bullet_{L,R},\circ_{c_{1}},\bullet_{c_{2}}\rangle\pm\lvert\bullet_{L,R},\bullet_{c_{1}},\circ_{c_{2}}\rangle)/\sqrt{2}$, and the matter basis set $\{|n_{L,R},n_{c_{1}},n_{c_{2}}\rangle\}$ has been used (number occupation representation of the non-local fermion $f_{L,R}$ and two local fermions fully contained inside the cavity, $c_{1}$ and $c_{2}$, see SM [28]). The result in Eq. (6) indicates that the system’s ground state is an entangled KC-cavity state: the symmetric combination of the two local fermion ($c_{1},c_{2}$) state couples with an even Schrödinger cat state whereas the anti-symmetric two fermion state couples to the odd Schrödinger cat state. Therefore, conditioned on the symmetry measurement outcome of the fermion state inside the $2CK$, an even or odd cavity cat state can be distilled. In the resonant case, the ground state of the Hamiltonian given by Eq. (2) is numerically obtained through exact diagonalization from which the reduced photon density matrix immediately is within reach for the GS out of the special point $\lambda_{1}=\lambda_{2}=\lambda$, $\phi_{1}=0$, and $\phi_{2}=\pi$. As depicted in the upper-lateral plots in Fig. 5 the Wigner functions do not display negative features as expected for coherent states. However, for the special set of coupling strengths we are considering in Eq. (6), after a previous projection of the GS on the matter states $\lvert\tilde{\Psi}_{\pm}\rangle$, the corresponding Wigner function displays negative values in the central plots of Fig. 5, a fingerprint of Majorana induced Schrödinger cat states.
Importantly, the generation of Schrödinger cat states in the $2CK$ system is independent of the resonance condition or dispersive condition, and neither a nonlinear Kerr medium is required. Furthermore, the dissipation effect of losing one excitation results, at some random time, in an inversion of parity of the cat state, i.e. even cat states transform to odd cat states (and vice versa).
Figure 5: Majorana-Schrödinger cat states. In the main panels (a) and (b), the bosonic Wigner function is represented, conditioned on the symmetry measurement outcome of the fermion state inside the 2CK in the resonance case. An even (panel (a)) or odd (panel (b)) cavity cat state is distilled. The red lines over the density projections correspond to the values of $\alpha=\pm\lambda/(\sqrt{2}\omega)$, with $\lambda=4$ fixed. Similarly, the insets correspond to the numerical Wigner functions obtained when the ground state is projected onto $\left|+,-\right\rangle_{x}$ and $\left|-,+\right\rangle_{x}$, respectively.
Summary and discussion.— In this work, we report on a diverse range of new features of anyons that can be significant for harnessing cavity confinement control over quantum topological materials. Our results not only uncover new facets of anyon-photon composed states at a fundamental level but also have interesting applications, such as braiding-allowed teleportation schemes in the USC regime. Our findings provide useful insights into how superconducting circuit cavity control is indeed a feasible alternative to test non-Abelian anyon statistics associated with MZMs. This would provide a stringent test of the theory of free Majorana fermions in nanowire or quantum dot array scenarios.
Enhanced experimental capabilities to control superconducting quantum electrodynamics circuits are currently fuelling intense research in the domain of light-matter systems. Harnessing ultra-strong couplings is thus not only of fundamental interest but also of appeal for technological applications. Regarding the experimental feasibility of the above-discussed results, it is worth noting that the KC model has been demonstrated experimentally in a fully controlled chain of coupled quantum dots where fine-tuning of parameters lays out the system into a sweet spot which leads to the generation of poor man’s Majorana states [6]. The next step should be the coupling to a transmission line resonator. With properly applied driving on the resonator mode or directly on the KC sites embedded within the cavity, the cavity-controlled KC features can be observed from the transmission spectrum [23, 24, 36, 37], by quantum process tomography [38] or by microwave interferometric measurements [39].
An interesting prospect is to generalize these single cavity controlled anyon processes to multi-cavity setups offering tools to observe geometric phase controlled braiding mechanisms offering new paths for further exploration. Besides, dissipation is something that cannot be avoided in experiments and thus should be accounted for in studying the photon-matter correlation transitions.
Acknowledgments.— L.Q. and F.J.R. would like to thank Facultad de Ciencias-UniAndes projects: INV-2021-128-2292 and INV-2023-162-2833 for financial support. F.J.G-R and I.A.B-G acknowledge financial support from Spanish MCIN with funding from European Union Next Generation EU (PRTRC17.I1) and Consejeria de Educacion from Junta de Castilla y Leon through QCAYLE project, as well as grants PID2020-113406GB-I00 MTM funded by AEI/10.13039/501100011033, and RED2022-134301-T. C.T. acknowledges financial support from the project “Nanofotonica para Computacion Cuantica” (NanoQuCo) with reference Y2020/TCS-6545 of the Comunidad de Madrid (CAM).